OpenStudy (anonymous):

what does it mean when the directions say write as a ratio of whole numbers in lowest terms???????????

OpenStudy (anonymous):

It means simplify the fraction. So the lowest term of 5/10 is 1/2

OpenStudy (anonymous):

oh ok how did u figure that out?

OpenStudy (anonymous):

If you mean the term "lowest term" i know because its just a synonym. If you mean how do I know its one half, you divide both numbers by something they have in common. The biggest number 10 and 5 have in common is 5. So 5/5 = 1 and 10/5 = 2

OpenStudy (anonymous):

ooo

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what if it says 1. 32 to 104

OpenStudy (anonymous):

same direction

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I don't think that one can be simplified. So the answer would be the same

OpenStudy (anonymous):

ohh how do u no that??

OpenStudy (anonymous):

the greatest common factor between 1.32 and 104 is 1. And if you divided both numbers by 1, you would still get the same two numbers. That's how I was taught but I think your directions might have to do with something else.

OpenStudy (anonymous):

ohhhhhh

OpenStudy (anonymous):

number 1

OpenStudy (anonymous):

Actually nevermind I thought you said 1.32 not 32 The greatest number you can divide both 32 and 104 by is 8 32/8 = 4 104/8 = 13 So your answer is 4/13

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

4 1/2 cups for 2 dozen

OpenStudy (anonymous):

watabout this?

OpenStudy (anonymous):

That one is a bit difficult because you have to convert either cups or dozens to the other one which I don't know a conversion factor for you might want to ask that in a separate question.